How do you solve for x?
ether Ko - Center water 10.0 = (X/100) (5.00g - X)/300
Answer #1

10.0 - (5.00g -X)/300 10.0 x(5.009-x) = مرد 10.0x (5.009-x) = 3x 50.00g - 10.0% = 3 50.oog E 10.0x + 3x = 13.0x 50.00g X = 50